A proton is fired from far away toward the nucleus of a mercury atom. Mercury is element number 80, and the diameter of the nucleus is 14.0 fm.
If the proton is fired at a speed of 4.6×107 m/s , what is its closest approach to the surface of the nucleus? Assume the nucleus remains at rest.
Answer #1

let r be the distance in closest approach

initial K.E = P.E

0.5 * mp * v^2 = (k *q1 * q2 / r)

(0.5 * 1.67 * 10^-27 * sqr(4.6* 10^7)) = (9*10^9 * 1.6*10^-19 * 80 * 1.6*10^-19 / r )

therefore r = 1.0432 * 10^-12 m